OpenStudy (zpupster):

A tennis ball bounces so that its initial speed straight upwards is b feet per second. Its height s in feet at time t seconds is given by s = bt − 16t2 a) Find the velocity v = ds/dt at time t. b) Find the time at which the height of the ball is at its maximum height. c) Find the maximum height.

OpenStudy (zpupster):

found a I just differentiate leaving b-32t let v(0) 0=b-32t t=b/32 finding the max height s(b/32) when velocity is 0 but they answer that---s( b/32)=b^2/64 why would they put b to a raised power. like they did in the answer.

OpenStudy (perl):

need help?

OpenStudy (perl):

s(b/32) = b(b/32) − 16(b/32)^2 = b^2 / 32 - 16 *b^2 / 1024 = 32b^2 / 1024 - 16b^2 / 1024 = 16b^2 / 1024 = b^2 / 64

OpenStudy (zpupster):

ty I plgged it in but received a different answer. I made a mistake somewhere. ty again